A. Solve the following.


1. If price increases from 25 to 35 (P=30) and quantity demanded falls from 80 to 60 units (Q=70), what is the elasticity of demand?


2. If price decreases from 40 to 30 and quantity demanded rises from 15 to 30 units, what is the elasticity of demand?


3. If income rises 15 percent in a year and the demand for clothing by 20 percent, find the YED.


4. If 5 percent increase in the of iphone, the demand for Huawei phone increases by 20 percent, find the XED.


1
Expert's answer
2021-10-02T10:48:05-0400

1. If price increases from 25 to 35 (P=30) and quantity demanded falls from 80 to 60 units (Q=70), then the elasticity of demand is:

"Ed = \frac{60-80} {35-25} ×\frac{35+25} {60+80} = -0.86."

So, the demand is inelastic.

2. If price decreases from 40 to 30 and quantity demanded rises from 15 to 30 units, then the elasticity of demand is:

"Ed = \frac{30-15} {30-40} ×\frac{30+40} {30+15} = -2.33."

So, the demand is elastic.

3. If income rises 15 percent in a year and the demand for clothing by 20 percent, then the YED is:

YED = 20/15 = 1.33.

So, clothing is a normal good.

4. If 5 percent increase in the price of iphone, the demand for Huawei phone increases by 20 percent, then the XED is:

XED = 20/5 = 4.

So, iphone and Huawei phone are substitutes.